OpenStudy (anonymous):

Find the antiderivative using Trigonometric Substitution: (cosx)^3

OpenStudy (anonymous):

rewrite (cosx)^3 = cos^2x 8 cosx = (1 - sin^2x)*cosx let u=sinx, du=cosx dx so your integral is now

OpenStudy (anonymous):

integral (1-u^2) du

OpenStudy (anonymous):

oh.. trig substitution? why?

OpenStudy (anonymous):

u-sub works fine here...